OpenStudy (christinasc4):

If the width of a rectangular field is 17 kilometers and the field has an area of 578 square kilometers what is the length of the field

jimthompson5910 (jim_thompson5910):

In this problem, the knowns are Width = W = 17 Area = A = 578 the unknown is Length = L = x Now we move onto the formula area of rectangle = length times width which is the same as A = L*W then plug in the knowns and solve for the unknown 578 = x*17 x = ??? I'll let you do the rest

OpenStudy (christinasc4):

Thank you. I needed to verify I was doing the formula correctly

jimthompson5910 (jim_thompson5910):

no problem